OpenStudy (anonymous):

HELP !! If the lines 4y+2x = -5 and 3y +ax = -2 are perpendicular, what is the value of a ?

OpenStudy (anonymous):

@Shido88 May I help you with this question?

OpenStudy (anonymous):

Of course !!

OpenStudy (anonymous):

Great! You must first transform the the equation 4y+2x = -5 into slope intercept form. Do you know how to do that?

OpenStudy (anonymous):

isn't it ax+b ?

OpenStudy (anonymous):

yes it is. so subtract 2x from the left side and divide all the terms by 4 to get y=ax+b

OpenStudy (anonymous):

ok hold on

OpenStudy (anonymous):

y= -2x -5/ 4 ?

OpenStudy (anonymous):

y= -1/2x - 5/4

OpenStudy (anonymous):

you forgot to divide -2x by four.

OpenStudy (anonymous):

oh i see

OpenStudy (anonymous):

got it! what next ?

OpenStudy (anonymous):

great! now that you have that equation, you know that the first slope is -1/2. do you follow? : )

OpenStudy (anonymous):

so what would be the slope of the perpendicular equation?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

2/1 ?

OpenStudy (anonymous):

yes or just plain 2 :) Give me one second and i will tell you the last part.

OpenStudy (anonymous):

sorry for the delay

OpenStudy (anonymous):

sure! take your time!

OpenStudy (anonymous):

my goodness. one second. i'm lost i thought i remembered doing this.

OpenStudy (anonymous):

aha it's ok, im here all day !

OpenStudy (anonymous):

haha all right im sorry about that

OpenStudy (anonymous):

^^

OpenStudy (anonymous):

i'm going to repost this question for you. i will talk you in the post.

OpenStudy (anonymous):

sure